# Aligning Model Properties via Conformal Risk Control

### v1

@@ 3 Conformal property alignment
Our main methodology is to use conformal risk control to create prediction intervals that align with specific properties $\mathcal{P}$. Our approach allows us to post-process the outputs of a pre-trained model $f$ to ensure that within the resulting conformal band, with a given probability, there exists predictions that adhere to desired properties such as monotonicity. We first present a generalization of conformal risk control (Angelopoulos et al., 2024) that exends to multi-dimensional $\boldsymbol{\lambda}$.
@@ 3.1 Multi-lambda conformal risk control
We generalize the conformal risk control algorithm to a $k$-dimensional vector of lambda values $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{k})$, where larger $\boldsymbol{\lambda}\in\mathbb{R}^{k}$ yield more conservative outputs. The construction of $C_{\boldsymbol{\lambda}}(X_{i})$ depends on the application and allows for flexibility in the use of $f(X_{i})$ and $\boldsymbol{\lambda}$.
Definition 2 ($C_{\boldsymbol{\lambda}}(X)$ and $C_{\boldsymbol{\lambda}}(f)$).
For each point $X_{i}\in\mathcal{X}$, define $C_{\boldsymbol{\lambda}}(X_{i})$ as the conformal set of predictions for $X_{i}$ constructed from $f$ and $\boldsymbol{\lambda}$. Then, $C_{\boldsymbol{\lambda}}(f)$ is the conformal interval around $f$ that includes $C_{\boldsymbol{\lambda}}(X_{i})$ for every $X_{i}\in\mathcal{X}$.
Following the original scalar $\lambda$ setting (Angelopoulos et al., 2024), we assess $C_{\boldsymbol{\lambda}}$ using non-increasing random loss functions $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i}),Y_{i})\in(-\infty,B]$ for $B<\infty$. In particular, we consider an exchangeable collection of non-increasing random functions $L_{i}:\boldsymbol{\Lambda}\to(-\infty,B],i=1,...,n+1$, where $\boldsymbol{\Lambda}\subset\mathbb{R}_{\geq 0}^{k}$ is a bounded subset of $\mathbb{R}_{\geq 0}^{k}$, with bound $\lambda^{\max}_{j}$ in each dimension $j\in[k]$.
@@
to be the set of minimal elements (Boyd and Vandenberghe, 2004) of $\boldsymbol{\Lambda}$ that satisfy the condition $\frac{n}{n+1}\hat{R}_{n}(\boldsymbol{\lambda})+\frac{B}{n+1}\leq\alpha$. Let $g:\boldsymbol{\Lambda}\to\mathbb{R}$ be a strictly increasing function with respect to the product order on $\mathbb{R}^{k}$. Thus each $L_{i}(\boldsymbol{\lambda})$ is non-increasing with respect to the level sets defined by $g(\boldsymbol{\lambda})$. Then select $\boldsymbol{\hat{\lambda}}\in\boldsymbol{\Lambda}_{\min}$ to be a minimizer of $g$ over $\boldsymbol{\Lambda}_{\min}$.
We then deploy the resulting conformal band $C_{\boldsymbol{\hat{\lambda}}}$ on the test point $X_{n+1}$. For this choice of $\boldsymbol{\hat{\lambda}}$, we have a risk control guarantee that mimics the result of Angelopoulos et al. (2024), specifically:
@@
Assume that $L_{i}(\boldsymbol{\lambda})$ is non-increasing with respect to the partial ordering of $\boldsymbol{\Lambda}$ inherited from the product order on $\mathbb{R}^{k}$. Additionally let $g(\boldsymbol{\lambda})$ be a strictly increasing function $g:\boldsymbol{\Lambda}\to\mathbb{R}$. Also assume $L_{i}$ is right-continuous in each dimension, $L_{i}(\boldsymbol{\lambda}^{\max})\leq\alpha$, and $\sup_{\boldsymbol{\lambda}}L_{i}(\boldsymbol{\lambda})\leq B<\infty$ almost surely. Then
@@
The proof is similar to the proof of the guarantee for the conformal risk control algorithm in Angelopoulos et al. (2024) and is deferred to Appendix A.
@@ 3.2 Conformal property alignment from proximity oblivious testers
Now we show how to construct a conformal risk control problem from proximity oblivious testers for a given property ${\mathcal{P}}$. Let $\mathsf{T}(f,X,Y)$ be the random output of $\mathsf{T}^{f,{\mathcal{D}}}$ evaluated at $(X,Y)\sim D$. Note there can be randomness in the machine besides sampling $(X,Y)\sim{\mathcal{D}}$, and each individual loss function will be defined with respect to fixing this randomness:
Definition 3 ($C_{\boldsymbol{\lambda}}(\mathsf{T})$).
We set $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i}),Y_{i})=0$ if after fixing the remaining randomness in $\mathsf{T}$, there exists $\tilde{f}$ such that $\mathsf{T}(\tilde{f},X_{i},Y_{i})=\text{Accept}$ where $\tilde{f}$ is the function resulting from modifying $f$ to take some value $\tilde{Y}\in C_{\boldsymbol{\lambda}}(X_{i})$ at $X_{i}$. Otherwise, $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i}),Y_{i})=1$. Define the set of loss functions obtained from this mapping to be $C_{\boldsymbol{\lambda}}(\mathsf{T})$.
Intuitively, the POT may fail for the original value of the function $f(X_{i})$ at $X_{i}$, but if we can adjust the value of $f$ to take some value $\tilde{Y}$ that falls within $C_{\boldsymbol{\lambda}}(X_{i})$ such that the function now satisfies the query at point $X_{i}$, then we take the loss to be zero. We use conformal risk control to find a $\boldsymbol{\lambda}$ such that with a sufficiently high probability this loss will be zero at a random data point. The goal is not to obtain a specific new function, but rather a band around $f$ for which we can establish probabilistic guarantees of predictions within the band satisfying ${\mathcal{P}}$ as determined by $\mathsf{T}$.
Linearity example.
@@
Note that in this case the tester and loss function does not depend on the $Y_{i}$. This is because the property of $f$ being monotone does not depend on the $Y_{i}$ and here ${\mathcal{D}}$ is only used to obtain samples of the $X_{i}$. This is not the case in general, however, and properties can be defined with respect to the whole sample $(X_{i},Y_{i})\sim{\mathcal{D}}$. For example, we could consider the property ${\mathcal{P}}$ that $f$ does not over-predict, that is, for $(X,Y)\sim{\mathcal{D}}$ we have $f(X)\leq Y$. Now we state our main theorem.
@@
If we run conformal risk control on a calibration dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from a distribution ${\mathcal{D}}$ using parameter $\alpha$ and loss functions $L=C_{\boldsymbol{\lambda}}(\mathsf{T})$ mapped from a proximity oblivious tester $\mathsf{T}$ for property ${\mathcal{P}}$, then with probability at least $1-\alpha$ there exists a function $\tilde{f}\in C_{\boldsymbol{\hat{\lambda}}}(f)$ such that $\tilde{f}$ is at most $\alpha$-far from ${\mathcal{P}}$ with respect to ${\mathcal{D}}$.
@@
Conformal risk control guarantees us that ${\mathbb{E}}[L_{n+1}(C_{\boldsymbol{\hat{\lambda}}}(X_{n+1}),Y_{n+1})]\leq\alpha$. Note that the randomness in this expectation is over the randomness in all of the $n+1$ points. We have
$\alpha$ $\geq{\mathbb{E}}[L_{n+1}(C_{\boldsymbol{\hat{\lambda}}}(X_{n+1}),Y_{n+1})]$
$\alpha$ $\geq\Pr_{(X_{1},Y_{1}),...,(X_{n+1},Y_{n+1})\sim{\mathcal{D}}}[\forall\tilde{f}\in C_{\boldsymbol{\hat{\lambda}}}(f),\mathsf{T}(\tilde{f},X_{n+1},Y_{n+1})=\text{Reject}]$
$1-\alpha$ $\leq\Pr_{(X_{1},Y_{1}),...,(X_{n+1},Y_{n+1})\sim{\mathcal{D}}}[\exists\tilde{f}\in C_{\boldsymbol{\hat{\lambda}}}(f),\mathsf{T}(\tilde{f},X_{n+1},Y_{n+1})=\text{Accept}]$
$1-\alpha$ $\leq\Pr_{(X_{n+1},Y_{n+1})\sim{\mathcal{D}}}[\exists\tilde{f}\in C_{\boldsymbol{\hat{\lambda}}}(f),\mathsf{T}(\tilde{f},(X_{n+1},Y_{n+1}))=\text{Accept}|\boldsymbol{\hat{\lambda}}]\Pr_{(X_{1},Y_{1}),...,(X_{n},Y_{n})\sim{\mathcal{D}}}[\boldsymbol{\hat{\lambda}}]$
So with probability at least $1-\alpha$ over the random draw of $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ we have that $\boldsymbol{\hat{\lambda}}$ is such that with probability at least $1-\alpha$ over sample $(X_{n+1},Y_{n+1})\sim{\mathcal{D}}$ we can find a point $\tilde{Y}_{n+1}\in C_{\boldsymbol{\hat{\lambda}}}(X_{n+1})$ such that the POT $\mathsf{T}$ accepts. For the points $(X,Y)$ in the support of ${\mathcal{D}}$ for which we can find such a $\tilde{Y}\in C_{\boldsymbol{\lambda}}(X)$, set $\tilde{f}(X)=\tilde{Y}$. Then define $\tilde{g}(X)$ to agree with $\tilde{f}(X)$ at these points and take a value such that $\mathsf{T}$ accepts at other points in ${\mathcal{X}}\times{\mathcal{Y}}$. Thus $\tilde{g}\in{\mathcal{P}}$. Then we have that $\Pr_{(X,Y)\sim{\mathcal{D}}}[\tilde{f}(X)\not=\tilde{g}(X)]\leq\alpha$. So $\min_{g\in{\mathcal{P}}}\Pr_{(X,Y)\sim D}[\tilde{f}(X)\not=g(X)]\leq\alpha$, hence $\tilde{f}$ is at most $\alpha$ far from ${\mathcal{P}}$ with respect to ${\mathcal{D}}$. ∎
@@ 4.1 Monotonicity
Consider a user given a pre-trained model $f$ that was not trained with monotonic constraints. The user, however, wishes for the sake of safe or fair deployment to make predictions in a way that is as monotonic as possible. In particular, let ${\mathcal{P}}$ be the property that $f$ is monotonically decreasing in dimension $k$. To apply our methodology we consider the following proximity oblivious tester $\mathsf{T}$ in Algorithm 1.
Algorithm 1 POT $\mathsf{T}$ for monotonicitally decreasing in a fixed dimension $k$
1:  Sample $(X_{1},Y_{1})\sim\mathcal{D}$. Let $X_{1}=(x_{1},x^{-k})$
2:  Sample $x_{2}$ uniformly from the range of dimension $k$. Set $X_{2}=(x_{2},x^{-k})$
3:  if $x_{1}<x_{2}$ and $f(X_{1})<f(X_{2})$ then
@@
Clearly this tester will always return Accept when $f\in{\mathcal{P}}$, and if $\delta_{{\mathcal{P}},{\mathcal{D}}}(f)>0$, then there will be some nonzero probability that $\mathsf{T}$ returns Reject. We apply our conformal property alignment from POT procedure to obtain loss functions $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i}),Y_{i})$. On our calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ we obtain a loss function $L_{i}$ for each $i$ that is obtained from fixing the randomness in $\mathsf{T}$ to obtain a $x_{i2}$ for each point $i$. Note that after fixing this randomness, since the calibration data point $(X_{i},Y_{i})$ is also fixed, each loss function $L_{i}$ becomes a deterministic function of $\boldsymbol{\lambda}$.
Now according to our construction of the loss functions in Section 3.2, we have that $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i1}),Y_{i})=0$ if there exists $\tilde{f}$ such that, with the fixed randomness of sampling $X_{i2}$, we have $\mathsf{T}(\tilde{f},X_{i1})=\text{Accept}$, where $\tilde{f}$ is the function resulting from modifying $f$ to take some value $\tilde{Y}\in[f(X_{i1})-\lambda^{-},f(X_{i1})+\lambda^{+}]$. Otherwise, $L_{i}=\ell(C_{\boldsymbol{\lambda}}(X_{i1}),Y_{i})=1$. Intuitively, we have zero loss if we can find some point in the conformal band around around $f(X_{i1})$ such that monotonicity (decreasing) is preserved with respect to $X_{i2}$ in the sense that monotonicity POT $\mathsf{T}$ returns Accept.
@@
We align for monotonicity on various UCI ML repository datasets (Dua and Graff, 2023) with a 70-15-15 train-calibrate-test split, averaged over 10 random splits. We use XGBoost regression models (Chen and Guestrin, 2016). For each dataset, we select a feature for which we desire the model to be monotonic, not with the intention of improving test-set accuracy, but from the perspective of a user who desires this property.

### v2

@@ 3 Conformal property alignment
Our main methodology is to use conformal risk control to create prediction intervals that align with specific properties $\mathcal{P}$. Our approach allows us to post-process the outputs of a pre-trained model $f$ to ensure that within the resulting conformal band, with a given probability, there exists predictions that adhere to desired properties such as monotonicity.
@@ 3.1 Multi-lambda conformal risk control
We make particular use of the conformal risk control algorithm to allow for a $k$-dimensional vector of tuning parameters $\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{k})$, where larger values of $\boldsymbol{\lambda}\in\boldsymbol{\Lambda}\subset\mathbb{R}^{k}$ yield more conservative outputs, where $\boldsymbol{\Lambda}\subset\mathbb{R}_{\geq 0}^{k}$ is a bounded subset of $\mathbb{R}_{\geq 0}^{k}$. This works by mapping $\boldsymbol{\lambda}$ to a scalar and then applying standard conformal risk control. We emphasize that this result is not new and follows essentially directly from Angelopoulos et al. (2024). The construction of the output set $F_{\boldsymbol{\lambda}}(X)\subseteq{\mathcal{Y}}$ depends on the specific application and provides flexibility in how the function $f(X)$ and the parameters $\boldsymbol{\lambda}$ are utilized.
Definition 5 (Construction of $F_{\boldsymbol{\lambda}}(X)$).
Let $f:\mathcal{X}\to\mathcal{Y}$ be a given function. For each $\boldsymbol{\lambda}\in\mathbb{R}^{k}$, define the set-valued function $F_{\boldsymbol{\lambda}}:\mathcal{X}\to 2^{\mathcal{Y}}$ such that, for each $X\in\mathcal{X}$, $F_{\boldsymbol{\lambda}}(X)$ is a set of predictions for $X$ constructed from $f$ and $\boldsymbol{\lambda}$. The specific construction of $F_{\boldsymbol{\lambda}}(X)$ should satisfy the following properties:
  1. 1.
    When $\boldsymbol{\lambda}=\boldsymbol{0}$, we have $F_{\boldsymbol{0}}(X)=\{f(X)\}$.
  2. 2.
    For any $\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}\in\mathbb{R}^{k}$, if $\boldsymbol{\lambda}\leq\boldsymbol{\lambda}^{\prime}$ (i.e., $\lambda_{i}\leq\lambda_{i}^{\prime}$ $\forall i=1,2,\ldots,k$), then $F_{\boldsymbol{\lambda}}(X)\subseteq F_{\boldsymbol{\lambda}^{\prime}}(X)$.
This definition ensures that increasing the parameters $\boldsymbol{\lambda}$ leads to larger (more conservative) prediction sets, and that when all parameters are zero, the prediction set reduces to the point prediction given by $f(X)$.
Following the original scalar $\lambda$ setting, we assess $F_{\boldsymbol{\lambda}}$ using non-increasing random loss functions $L_{i}=\ell(F_{\boldsymbol{\lambda}}(X_{i}),Y_{i})\in(-\infty,B]$ for $B<\infty$. In particular, we consider an exchangeable collection of non-increasing random functions $L_{i}:\boldsymbol{\Lambda}\to(-\infty,B],i=1,...,n+1$, where $\boldsymbol{\Lambda}\subset\mathbb{R}_{\geq 0}^{k}$ is a bounded subset of $\mathbb{R}_{\geq 0}^{k}$, with bound $\lambda^{\max}_{j}$ in each dimension $j\in[k]$.
@@
to be the set of minimal elements (Boyd and Vandenberghe, 2004) of $\boldsymbol{\Lambda}$ that satisfy the condition $\frac{n}{n+1}\hat{R}_{n}(\boldsymbol{\lambda})+\frac{B}{n+1}\leq\alpha$. Let $g:\boldsymbol{\Lambda}\to\mathbb{R}$ be a strictly increasing function such that $L_{i}(\boldsymbol{\lambda})$ is non-increasing with respect to the level sets defined by $g(\boldsymbol{\lambda})$. Then select $\boldsymbol{\hat{\lambda}}\in\boldsymbol{\Lambda}_{\min}$ to be a minimizer of $g$ over $\boldsymbol{\Lambda}_{\min}$.
We then deploy the resulting set-valued function $F_{\boldsymbol{\hat{\lambda}}}$ on the test point $X_{n+1}$. For this choice of $\boldsymbol{\hat{\lambda}}$, we have a risk control guarantee that mimics the result of Angelopoulos et al. (2024), specifically:
@@
Assume that $L_{i}(\boldsymbol{\lambda})$ is non-increasing with respect to the partial ordering of $\boldsymbol{\Lambda}$ inherited from $\mathbb{R}^{k}$. Additionally, assume that $L_{i}(\boldsymbol{\lambda})$ is non-increasing with respect to $g(\boldsymbol{\lambda})$ for some strictly increasing function $g:\boldsymbol{\Lambda}\to\mathbb{R}$. Also assume $L_{i}$ is right-continuous in each dimension, $L_{i}(\boldsymbol{\lambda}^{\max})\leq\alpha$, and $\sup_{\boldsymbol{\lambda}}L_{i}(\boldsymbol{\lambda})\leq B<\infty$ almost surely. Then
@@
The proof is similar to the proof of the guarantee for the conformal risk control algorithm in Angelopoulos et al. (2024) and is deferred to Appendix C.
@@ 3.2 Conformal property alignment from proximity oblivious testers
We now demonstrate how to construct a conformal risk control problem using proximity-oblivious testers (POTs) for a given property $\mathcal{P}$. Suppose we are given a pre-trained model $f:\mathcal{X}\to\mathcal{Y}$. We aim to extend the point predictions of $f$ to prediction sets, where the size or conservativeness of the set is parameterized by a parameter $\boldsymbol{\lambda}$. Let $F_{\boldsymbol{\lambda}}:\mathcal{X}\to 2^{\mathcal{Y}}$ denote the set-valued function that outputs, for each $X\in\mathcal{X}$, the set $F_{\boldsymbol{\lambda}}(X)\subseteq\mathcal{Y}$ determined by $f$, $X$, and $\boldsymbol{\lambda}$.
Let $\mathcal{T}^{F,\mathcal{D}}$ be a proximity-oblivious tester for whether a set-valued function $F$ accommodates the property ${\mathcal{P}}$ as given y Definition 4. We denote the random output of $\mathcal{T}^{F,\mathcal{D}}$ evaluated at $(X,Y)\sim\mathcal{D}$ by $\mathcal{T}^{F,\mathcal{D}}(X,Y)$.
We now define a loss function, generated from $\mathcal{T}^{F,\mathcal{D}}$, which will be crucial in formulating our conformal risk control problem.
Definition 6 (Loss Function Generated from a POT).
Let $\mathcal{T}^{F,\mathcal{D}}$ be a proximity-oblivious tester for a property $\mathcal{P}$. We define the loss function $L_{i}$ as:
$L_{i}=\begin{cases}0,&\text{if }\mathcal{T}^{F,\mathcal{D}}(X_{i},Y_{i})=\text{Accept},\\ 1,&\text{otherwise},\end{cases}$
where $(X_{i},Y_{i})$ are samples from the distribution $\mathcal{D}$.
Example.
@@
$F_{\boldsymbol{\lambda}}(X)=[f(X)-\lambda^{-},\ f(X)+\lambda^{+}],$
for each $X\in\mathbb{R}$. This creates prediction intervals around the point prediction $f(X)$, with widths controlled by $\lambda^{-}$ and $\lambda^{+}$.
We then apply the loss function generated by $\mathcal{T}^{F_{\lambda},\mathcal{D}}$ as given in Definition 6, and use conformal risk control to tune $\boldsymbol{\lambda}$ such that the expected loss on the $(n+1)$th point falls below a given target level $\alpha$.
Note that in this case the tester and loss function does not depend on the $Y_{i}$. This is because the property of $f$ being constant does not depend on the $Y_{i}$ from the calibration set and here ${\mathcal{D}}$ is only used to obtain samples of the $X_{i}$. This is not the case in general, however, and properties can be defined with respect to the whole sample $(X_{i},Y_{i})\sim{\mathcal{D}}$. For example, we could consider the property ${\mathcal{P}}$ that $f$ does not over-predict, that is, for $(X,Y)\sim{\mathcal{D}}$ we have $f(X)\leq Y$. Now we state our main theorem.
@@
Let $\mathcal{T}$ be a proximity-oblivious tester for a property $\mathcal{P}$ with detection probability function $\rho(\cdot)$. Assume access to a calibration dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$ sampled independently from a distribution $\mathcal{D}$. Suppose we run conformal risk control on this calibration dataset using risk parameter $\alpha$ and loss functions $L_{i}$ for property ${\mathcal{P}}$ generated from ${\mathcal{T}}$ (as in Definition 6). Then, for any $\varepsilon$ such that $\rho(\varepsilon)>\alpha$, the probability that $F_{\hat{\boldsymbol{\lambda}}}$ is $\varepsilon$-Faraway from $\mathcal{P}$ satisfies:
$\Pr_{(X_{1},Y_{1}),...,(X_{n},Y_{n})}\left(F_{\hat{\boldsymbol{\lambda}}}\text{ is }\varepsilon\text{-Faraway from }\mathcal{P}\right)\leq\frac{\alpha}{\rho(\varepsilon)}.$
@@
Let $\mathcal{E}$ denote the event that $F_{\hat{\boldsymbol{\lambda}}}$ is $\varepsilon$-Faraway from the property $\mathcal{P}$. Our goal is to bound the probability $\Pr_{(X_{1},Y_{1}),...,(X_{n},Y_{n})}[\mathcal{E}]$.
The conformal risk control procedure ensures that the expected loss on a new sample $(X_{n+1},Y_{n+1})$ satisfies:
$\mathbb{E}_{(X_{1},Y_{1}),...,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})}[L_{n+1}]\leq\alpha.$
Now, we can write
$\mathbb{E}[L_{n+1}]$ $=\Pr(\mathcal{E})\cdot\mathbb{E}[L_{n+1}\mid\mathcal{E}]+\Pr(\mathcal{E}^{c})\cdot\mathbb{E}[L_{n+1}\mid\mathcal{E}^{c}].$
When $\mathcal{E}$ occurs, $F_{\hat{\boldsymbol{\lambda}}}$ is $\varepsilon$-Faraway from $\mathcal{P}$. By the properties of the proximity-oblivious tester $\mathcal{T}$, we have:
$\Pr_{(X,Y)\sim\mathcal{D}}\left[\mathcal{T}^{F_{\hat{\boldsymbol{\lambda}}},\mathcal{D}}(X,Y)=\text{Reject}\mid\mathcal{E}\right]\geq\rho(\varepsilon).$
Thus, the conditional expected loss satisfies:
$\mathbb{E}[L_{n+1}\mid\mathcal{E}]=\Pr\left[\mathcal{T}^{F_{\hat{\boldsymbol{\lambda}}},\mathcal{D}}(X_{n+1},Y_{n+1})=\text{Reject}\mid\mathcal{E}\right]\geq\rho(\varepsilon).$
And when since $\Pr(\mathcal{E}^{c})\cdot\mathbb{E}[L_{n+1}\mid\mathcal{E}^{c}]$ is non-negative because $L_{n+1}$ is non-negative, we obtain
$\mathbb{E}[L_{n+1}]$ $\geq\Pr(\mathcal{E})\cdot\rho(\varepsilon).$
Combining this result with our guarantee from the conformal risk control procedure,
$\alpha\geq\mathbb{E}[L_{n+1}]\geq\Pr(\mathcal{E})\cdot\rho(\varepsilon).$
This implies:
$\Pr(\mathcal{E})\leq\frac{\alpha}{\rho(\varepsilon)}.$
Therefore, the probability that $F_{\hat{\boldsymbol{\lambda}}}$ is $\varepsilon$-Faraway from $\mathcal{P}$ satisfies:
$\Pr_{(X_{1},Y_{1}),...,(X_{n},Y_{n})}\left(F_{\hat{\boldsymbol{\lambda}}}\text{ is }\varepsilon\text{-Faraway from }\mathcal{P}\right)\leq\frac{\alpha}{\rho(\varepsilon)}.$
Amplifying Detection Probability via Independent Calls.
When the detection probability $\rho(\varepsilon)$ of the proximity-oblivious tester $\mathcal{T}$ is less than or close to the risk parameter $\alpha$, the bound provided by Theorem 1 may not be tight or meaningful (since $\alpha/\rho(\varepsilon)$ could be greater than or equal to 1). To address this issue, we can amplify the detection probability by performing multiple independent executions of $\mathcal{T}$ and combining their results appropriately.
To increase the detection probability beyond $\alpha$, we execute the proximity-oblivious tester $\mathcal{T}$ independently $k$ times on independent samples and define a new tester $\mathcal{T}^{\prime}$ that rejects if any of the $k$ executions reject (i.e., by applying a logical OR to the outcomes). This amplification technique yields an adjusted detection probability
$\rho^{\prime}(\varepsilon)=1-(1-\rho(\varepsilon))^{k},$
representing the probability that at least one of the $k$ independent executions rejects when the function is $\varepsilon$-Faraway from $\mathcal{P}$.
In this approach, the calibration dataset needs to be partitioned into $n^{\prime}=\left\lfloor\frac{n}{k}\right\rfloor$ disjoint batches, each containing $k$ samples. Each batch provides the independent samples required for the $k$ executions of $\mathcal{T}$ per calibration point. As a result, the effective sample size available for calibrartion becomes $n^{\prime}$ due to this batching of samples.
@@ 4.1 Monotonicity
Consider a user given a pre-trained model $f$ that was not trained with monotonic constraints. The user, however, wishes for the sake of safe or fair deployment to make predictions in a way that is as monotonic as possible. In particular, let ${\mathcal{P}}$ be the property that $f$ is monotonically decreasing in dimension $k$. To apply our methodology we consider the proximity oblivious tester ${\mathcal{T}}$ for ${\mathcal{P}}$ as given in Algorithm 1.
Algorithm 1 POT ${\mathcal{T}}$ for property ${\mathcal{P}}$ of monotonically decreasing in dimension $k$
1:  Sample $X_{1}\sim\mathcal{D}$. Let $X_{1}=(x_{1},x^{-k})$
2:  Sample $x_{2}$ from the marginal distribution of ${\mathcal{D}}$ in dimension $k$. Set $X_{2}=(x_{2},x^{-k})$
3:  if $x_{1}<x_{2}$ and $\max{F(X_{1})}<\min F(X_{2})$ then
@@
[…]
@@
We align for monotonicity on various UCI ML repository datasets (Dua and Graff, 2023) with a 70-15-15 train-calibrate-test split, averaged over 30 random splits. We use XGBoost regression models (Chen and Guestrin, 2016). For each dataset, we select a feature for which we desire the model to be monotonic, not with the intention of improving test-set accuracy, but from the perspective of a user who desires this property.